OpenStudy (anonymous):

substitution & addition 5x+y=-3 y=x-9

OpenStudy (saifoo.khan):

x = 1 y = -8

OpenStudy (anonymous):

thank you

OpenStudy (saifoo.khan):

you welcome

OpenStudy (anonymous):

5x+y=-3, so y=-5x-3, so to get the value for y we need to make x=0, so y=-5(0)-3 so y=-8 now we do the same to get the value for x, we make y=0 so 0=-5x-3, so 5x=-3 then x=-3/5 correct me if am wrong...